Insert the following IP addresses into a hash table of size 5 using universal hashing with a1 = 1, a2 = 2, a3 = 4 and a4 = 5.:

209.85.231.104 Google

207.46.170.123 Microsoft

208.80.152.2 Wikipedia
Answer #1

Ans.)Each IP address is a 32 bit, 4 tuple address as <x1,x2,x3,x4> where xi={0,.....,255}

therefore, hashing will be as:

ha=(x1,x2,x3,x4)= a1x1+a2x2+a3x3+a4x4

Data hashed value
209.85.231.104 209 x 1+85 x 2+231 x 4+ 104 x 5=1823
207.46.170.123 207 x 1+ 46 x 2+170 x 4+123 x 5=1594
208.80.152.2 208 x 1+80 x 2+152 x 4+2 x 5=986